# Augmenting Subspace Optimization Methods with Linear Bandits

### v1

@@ 3 Linear UCB Method for Playing Framework 2
For ease of analysis, we consider an extreme case where the environment in Framework 2 always chooses $p_{k}=0$, so that $S_{k}$ is empty; it is transparent how one would adapt the algorithm and analysis for the more general case, but bookkeeping and notation become tedious. In a practical sense, the environment always choosing $p_{k}=0$ is tantamount to performing absolutely no sketching via $S_{k}$. This trivial selection of $\{p_{k}\}$ better highlights the fact that the regret bounds we prove concern the satisfaction of (2) and are not actually related to the dynamics of the Framework 1 optimization method employed in Framework 2. It is also favorable from a practical point of view that the restriction of $p_{k}=0$ yields a deterministic method.
Pseudocode for our linear UCB method is provided in Algorithm 3.
@@
2(Inputs) Initial point $x_{1}$, horizon $K$, regularizer $\lambda>0$.
3(Initializations) Initial covariance matrix $C_{1}(\leftarrow\lambda I_{d})$, initial right-hand side $b_{1}(\leftarrow 0_{d})$, initial gradient estimate $g_{1}(\leftarrow 0_{d})$, stepsizes $\{\alpha_{k}\}>0$, memory parameter $M\geq 0$.
@@
6      Determine some upper bound $U_{k}\geq\|\nabla f(x_{k})\|$.
@@ 7      Compute
$s_{k}\leftarrow\displaystyle\arg\max_{s:\|s\|=1}g_{k}^{\top}s+\sqrt{\lambda}U_{k}\|s\|_{C_{k}^{-1}}$ (6)
@@
$r_{k}\leftarrow\langle\nabla f(x_{k}),s_{k}\rangle.$ (7)
@@
10      $C_{k+1}\leftarrow\lambda I_{d}+\displaystyle\sum_{j=\max\{1,k-M\}}^{k}s_{j}s_{j}^{\top}.$
11      $b_{k+1}\leftarrow\displaystyle\sum_{j=\max\{1,k-M\}}^{k}r_{j}s_{j}$
@@
$g_{k+1}\leftarrow C_{k+1}^{-1}b_{k+1}$ (8)
13      (Update incumbent.) $x_{k+1}\leftarrow x_{k}-\alpha_{k}r_{k}s_{k}$.
@@
Algorithm 3, and much of our analysis, is based on prior work in [14, 44, 59, 15]. As previously suggested, Algorithm 3 dynamically estimates $\nabla f(x_{k})$ via $g_{k}$ and moreover coarsely quantifies the decision-maker’s present uncertainty about $s_{k}^{\top}\nabla f(x_{k})$ in recently unexplored directions $s_{k}$ via the regularized covariance matrix $C_{k}$. This coarse uncertainty quantification is performed by regularized linear least squares, using the actions $s_{k}$ as linear measurements (maintained in the matrix $C_{k}$) and the directional derivatives $\langle\nabla f(x_{k}),s_{k}\rangle$ as responses (maintained in the vector $b_{k}$). We note the presence of a memory parameter $M$, which allows sufficiently old measurement/response pairs to be forgotten; this $M$ defines a so-called sliding window mechanism. Such forgetting is critical in a nonstationary bandit problem like that described in Framework 2. Together, $g_{k}$ and $C_{k+1}^{-1}$ define the isocontours of a confidence ellipse. The width of the confidence ellipse is parameterized by an estimated upper bound $U_{k}$ on $\|\nabla f(x_{k})\|$. For simplicity, we assume knowledge of such an upper bound that is valid for all $k=1,2,\dots,K$.
Assumption 4.
The gradient is bounded like $\|\nabla f(x)\|\leq G$ for some $G>0$ for all $x\in L(x_{0})$.
With 4, one could take $U_{k}=G$ for all $k$. However, because Algorithm 3 iteratively selects $s_{k}$ as a maximizer over the confidence ellipse in Equation 6, one can imagine that tighter upper bounds $U_{k}$ are generally preferable, so as to encourage better exploitation, that is, better identification of maximizers closer to the mean estimate $g_{k}$. We will discuss a practical means of dynamically adjusting $U_{k}$ in our numerical experiments. Note, however, that our analysis requires only an upper bound, and so 4 suffices for that purpose.
We set out to prove sublinear regret of Algorithm 3, but we first require two lemmata, the proofs of which are left to the Appendix.
@@
For each $k=1,2,\dots,K$, we have for any $s\in\mathbb{R}^{d}$ satisfying $\|s\|=1$,
$\|s^{\top}(\nabla f(x_{k})-g_{k})\|\leq\displaystyle\sqrt{\frac{d(M+1)}{\lambda}}\displaystyle\sum_{i=\ell(k)}^{k-1}\|\nabla f(x_{i})-\nabla f(x_{i+1})\|+\sqrt{\lambda}\|\nabla f(x_{k})\|\|s\|_{C_{k}^{-1}},$
where $\ell(k):=\max\{1,k-M-1\}$.
@@
Let $M\geq 1$, and denote $\ell(j)=\max\{1,j-M-1\}$. Let $C_{0}=\lambda I_{d}$, and let $C_{j}=C_{j-1}+s_{j}s_{j}-s_{\ell(j)}s_{\ell(j)}^{\top}$ for an arbitrary sequence of unit vectors $\{s_{j}\}_{j=1}^{K}\subset\mathbb{R}^{d}$. Then,
$\sum_{j=1}^{k}\|s_{j}\|_{C_{j-1}^{-1}}^{2}\leq\displaystyle\frac{2kd}{M}\log\left(1+\displaystyle\frac{M}{\lambda d}\right).$
We can now prove our main theorem. To make the statement precise, we define one additional piece of notation, the total variation of the sequence of gradients: that is,
$V_{K}:=\displaystyle\sum_{k=1}^{K}\|\nabla f(x_{k+1})-\nabla f(x_{k})\|.$
@@
Suppose $f$ satisfies 1. Given a sequence of gradients $\{\nabla f(x_{k})\}$, abbreviate $s_{k}^{*}:=\nabla f(x_{k})/\|\nabla f(x)\|.$ The cumulative dynamic regret of Algorithm 3, with $M\geq 1$, after $K$ iterations is bounded as
$D_{K}\leq\displaystyle\sum_{k=1}^{K}(s_{k}^{*}-s_{k})^{\top}\nabla f(x_{k})\leq 2M\displaystyle\sqrt{\frac{d(M+1)}{\lambda}}V_{K}+\displaystyle\sqrt{\frac{8\lambda d\sum_{k=1}^{K}U_{k}^{2}}{M}\log\left(1+\displaystyle\frac{M}{\lambda d}\right)}\sqrt{K}.$
@@
$\beta_{k}:=\displaystyle\sqrt{\frac{d(M+1)}{\lambda}}\displaystyle\sum_{i=\ell(k)}^{k-1}\|\nabla f(x_{i+1})-\nabla f(x_{i})\|.$
@@
$s_{k}^{*\top}\nabla f(x_{k})\leq s_{k}^{*\top}g_{k}+\beta_{k}+\sqrt{\lambda}\|\nabla f(x_{k})\|\|s_{k}^{*}\|_{C_{k}^{-1}}$
@@
$s_{k}^{\top}\nabla f(x_{k})\geq s_{k}^{\top}g_{k}-\beta_{k}-\sqrt{\lambda}\|\nabla f(x_{k})\|\|s_{k}\|_{C_{k}^{-1}}.$
@@
$\begin{array}[]{rl}(s_{k}^{*}-s_{k})^{\top}\nabla f(x_{k})&\leq(s_{k}^{*}-s_{k})^{\top}g_{k}+2\beta_{k}+\sqrt{\lambda}\|\nabla f(x_{k})\|\left[\|s_{k}^{*}\|_{C_{k}^{-1}}+\|s_{k}\|_{C_{k}^{-1}}\right]\\ &\leq(s_{k}^{*}-s_{k})^{\top}g_{k}+2\beta_{k}+\sqrt{\lambda}U_{k}\left[\|s_{k}^{*}\|_{C_{k}^{-1}}+\|s_{k}\|_{C_{k}^{-1}}\right]\\ &\leq 2\beta_{k}+2\sqrt{\lambda}U_{k}\|s_{k}\|_{C_{k}^{-1}},\end{array}$
where the last inequality is because $s_{k}$ is a maximizer in eq. 6. Thus,
$\begin{array}[]{rl}\displaystyle\sum_{k=1}^{K}(s_{k}^{*}-s_{k})^{\top}\nabla f(x_{k})&\leq 2\displaystyle\sum_{k=1}^{K}\beta_{k}+2\sqrt{\lambda}\displaystyle\sum_{k=1}^{K}U_{k}\|s_{k}\|_{C_{k}^{-1}}\\ &=2\displaystyle\sqrt{\frac{d(M+1)}{\lambda}}\sum_{k=1}^{K}\sum_{i=\ell(k)}^{k-1}\|\nabla f(x_{i+1})-\nabla f(x_{i})\|+2\sqrt{\lambda}\displaystyle\sum_{k=1}^{K}U_{k}\|s_{k}\|_{C_{k}^{-1}}\\ &\leq 2M\displaystyle\sqrt{\frac{d(M+1)}{\lambda}}V_{K}+2\sqrt{\lambda}\displaystyle\sum_{k=1}^{K}U_{k}\|s_{k}\|_{C_{k}^{-1}}\\ &\leq 2M\displaystyle\sqrt{\frac{d(M+1)}{\lambda}}V_{K}+2\sqrt{\lambda}\displaystyle\sqrt{\frac{2Kd}{M}\log\left(1+\displaystyle\frac{M}{\lambda d}\right)}\sqrt{\sum_{k=1}^{K}U_{k}^{2}},\\ \end{array}$
where we have used Lemma 4 and the Cauchy–Schwartz inequality to derive the last inequality. ∎
@@ 3.1 Understanding Theorem 1
To better understand Theorem 1, we will focus on the simple iteration (3). Suppose $\alpha_{k}=\frac{1}{L}$ for all $k$. We first note that, as a very loose bound,
$\|\nabla f(x_{k+1})-\nabla f(x_{k})\|\leq L\|x_{k+1}-x_{k}\|=L\|\frac{1}{L}P_{k}\nabla f(x_{k})\|\leq\|\nabla f(x_{k})\|$
@@
$V_{K}\leq\displaystyle\sum_{k=1}^{K}\|\nabla f(x_{k})\|.$ (9)
We stress that eq. 9 is practically quite loose; for many functions satisfying 1, the bound $\|\nabla f(x_{k+1})-\nabla f(x_{k})\|\leq L\|x_{k+1}-x_{k}\|$ is not usually tight, given that $L$ is a global Lipschitz constant per 1. Nonetheless, we will use the coarse upper bound in eq. 9 to derive Corollary 1 and Corollary 2. Moreover, we can further upper bound the right-hand side in (9) by using a straightforward result.
@@
Let $K>0$, and let $\delta^{^{\prime}}>0$. Let 2 hold, choosing $\delta=\delta^{^{\prime}}/K$, and letting $q>0$ be arbitrary. Let 3 hold. Then, the iterative method defined by eq. 3 satisfies, with probability at least $1-\delta^{^{\prime}}$,
$\sum_{k=0}^{K-1}\|\nabla f(x_{k})\|^{2}\leq\displaystyle\frac{2S_{\max}L(f(x_{0})-f(x_{*}))}{q}.$ (10)
@@
$\sum_{k=0}^{K-1}\|\nabla f(x_{k})\|\leq\displaystyle\sqrt{\displaystyle\frac{2S_{\max}L(f(x_{0})-f(x_{*}))}{q}}\sqrt{K}.$ (11)
@@
By Taylor’s theorem, we have for all $k\leq K-1$
$\begin{array}[]{rl}f(x_{k+1})&\leq f(x_{k})-\frac{1}{L}\nabla f(x_{k})^{\top}S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}\nabla f(x_{k})+\frac{1}{2L}\|S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}\nabla f(x_{k})\|^{2}\\ &=f(x_{k})-\frac{1}{L}\|S_{k}^{\top}\nabla f(x_{k})\|^{2}_{(S_{k}^{\top}S_{k})^{-1}}+\frac{1}{2L}\|S_{k}^{\top}\nabla f(x_{k})\|^{2}_{(S_{k}^{\top}S_{k})^{-1}}\\ &=f(x_{k})-\frac{1}{2L}\|S_{k}^{\top}\nabla f(x_{k})\|^{2}_{(S_{k}^{\top}S_{k})^{-1}}\\ &\leq f(x_{k})-\frac{1}{2S_{\max}L}\|S_{k}^{\top}\nabla f(x_{k})\|^{2},\\ \end{array}$
where 3 was used to derive the last inequality. By 2, with probability $1-(\delta^{^{\prime}}/K)$,
$f(x_{k+1})\leq f(x_{k})-\frac{q}{2S_{\max}L}\|\nabla f(x_{k})\|^{2}$
@@
$f(x_{0})-f(x_{K})\geq f(x_{0})-f(x_{K})\geq\displaystyle\frac{q}{2S_{\max}L}\displaystyle\sum_{k=0}^{K-1}\|\nabla f(x_{k})\|^{2}\quad\text{w.p. }1-\delta^{^{\prime}}.$
The use of an union bound in the proof of Theorem 2 is admittedly brutish. As a result, for very large horizons $K$, we require the probability $1-(\delta^{^{\prime}}/K)$ in 2 to be so close to 1 that regardless of how $S_{k}$ is generated, $p_{k}\approx d$. Moreover, Theorem 2 makes sense only when $K$ is finite, which is reasonable for our setting but would be generally nonsensical for performing rate analyses. One route to conducting a more refined analysis is to consider imposing a globalization strategy so that the high-probability convergence rate analysis of [9] could be imported here. Attempting this analysis would complicate our presentation, however, and distract from the purpose of this manuscript. Instead, we interpret Theorem 2 as telling us that, so long as an optimization algorithm used within Framework 2 yields similar first-order optimality guarantees to subspace gradient descent method in eq. 3, an upper bound like $V_{k}\in\mathcal{O}(\sqrt{K})$ is plausible. We can now state our two corollaries.
@@
Let 4 hold. Let the setting and assumptions of Theorem 2 hold. Suppose the iteration (3) is employed as the optimization algorithm in Framework 2 with a constant stepsize $\alpha_{k}=\frac{1}{L}$. Then, with probability at least $1-\delta^{^{\prime}}$,
$D_{K}\leq\left(2M\displaystyle\sqrt{\frac{2d(M+1)S_{\max}L(f(x_{0})-f(x_{*}))}{q\lambda}}+\displaystyle\sqrt{\frac{8\lambda dKG^{2}}{M}\log\left(1+\displaystyle\frac{M}{\lambda d}\right)}\right)\sqrt{K},$
and so if we choose $M=\lceil K^{\frac{1}{4}}\rceil$, then
@@
Additionally, we can prove a result for the idealized case where $U_{k}=\|\nabla f(x_{k})\|$ (that is, the upper bound used in eq. 6 is maximally tight) on every iteration.

### v4

@@ 3 Linear UCB Method for Playing Framework 2
For ease of analysis, we consider an extreme case where the environment in Framework 2 always chooses $p_{k}=0$, so that $\mathbf{S}_{k}$ is empty; it is transparent how one would adapt the algorithm and analysis for the more general case, but bookkeeping and notation become tedious. In a practical sense, the environment always choosing $p_{k}=0$ is tantamount to performing absolutely no sketching via $\mathbf{S}_{k}$. This trivial selection of $\{p_{k}\}$ better highlights the fact that the regret bounds we prove concern the satisfaction of (2) and are not actually related to the dynamics of the Framework 1 optimization method employed in Framework 2. It is also favorable from a practical point of view that the restriction of $p_{k}=0$ yields a deterministic method. As such, we make the critical observation that the results proven in Lemma 3, Lemma 4 and Theorem 1 are all deterministic results given a (trivial) sequence of $\{\mathbf{S}_{k}\}$.
We additionally comment that the linear UCB algorithm that we employ for this nonstationary environment most naturally works for one-dimensional sketches, since one-dimensional sketches correspond to the notion of having linearly parameterized actions. In future work, it may be interesting to attempt to solve a version of the subproblem (5) that optimizes over the Stiefel manifold of orthonormal $p_{k}$-frames for $p_{k}>1$, but this poses immediate questions about tractability.
Pseudocode for our linear UCB method is provided in Algorithm 4. We will note before stating Algorithm 4 that every iteration will require some upper bound $U_{k}\geq\|\nabla f(\mathbf{x}_{k})\|$. As we will demonstrate later in Corollary 1, simply knowing a global upper bound on the gradient norm is sufficient, i.e. we suppose we have knowledge of $G$ in the following assumption.
Assumption 4.
The gradient is bounded like $\|\nabla f(\mathbf{x})\|\leq G$ for some $G>0$ for all $\mathbf{x}\in L(\mathbf{x}_{0})$.
@@
2(Inputs) Initial point $\mathbf{x}_{1}$, horizon $K$, regularizer $\lambda>0$.
3(Initializations) Initial covariance matrix $\mathbf{C}_{1}(\leftarrow\lambda\mathbf{I}_{d})$, initial right-hand side $\mathbf{b}_{1}(\leftarrow\mathbf{0}_{d})$, initial gradient estimate $\mathbf{g}_{1}(\leftarrow\mathbf{0}_{d})$, stepsizes $\{\alpha_{k}\}>0$, memory parameter $M\geq 0$.
@@
6  Determine some upper bound $U_{k}\geq\|\nabla f(\mathbf{x}_{k})\|$.
@@ 7  Compute
$\mathbf{s}_{k}\leftarrow\displaystyle\arg\max_{\mathbf{s}:\|\mathbf{s}\|=1}\mathbf{g}_{k}^{\top}\mathbf{s}+\sqrt{\lambda}U_{k}\|\mathbf{s}\|_{\mathbf{C}_{k}^{-1}}$ (5)
@@
$r_{k}\leftarrow\langle\nabla f(\mathbf{x}_{k}),\mathbf{s}_{k}\rangle.$ (6)
@@
10 $\mathbf{C}_{k+1}\leftarrow\lambda\mathbf{I}_{d}+\displaystyle\sum_{j=\max\{1,k-M\}}^{k}\mathbf{s}_{j}\mathbf{s}_{j}^{\top}.$
11 $\mathbf{b}_{k+1}\leftarrow\displaystyle\sum_{j=\max\{1,k-M\}}^{k}r_{j}\mathbf{s}_{j}$
@@
$\mathbf{g}_{k+1}\leftarrow\mathbf{C}_{k+1}^{-1}\mathbf{b}_{k+1}$ (7)
13 (Update incumbent.) $\mathbf{x}_{k+1}\leftarrow\mathbf{x}_{k}-\alpha_{k}r_{k}\mathbf{s}_{k}$.
@@
Algorithm 4, and much of our analysis, is based on prior work in [15, 51, 67, 16]. As previously suggested, Algorithm 4 dynamically estimates $\nabla f(\mathbf{x}_{k})$ via $\mathbf{g}_{k}$ and moreover coarsely quantifies the decision-maker’s present uncertainty about $\mathbf{s}_{k}^{\top}\nabla f(\mathbf{x}_{k})$ in recently unexplored directions $\mathbf{s}_{k}$ via the regularized covariance matrix $\mathbf{C}_{k}$. This coarse uncertainty quantification is performed by regularized linear least squares, using the actions $\mathbf{s}_{k}$ as linear measurements (maintained in the matrix $\mathbf{C}_{k}$) and the directional derivatives $\langle\nabla f(\mathbf{x}_{k}),\mathbf{s}_{k}\rangle$ as responses (maintained in the vector $\mathbf{b}_{k}$). We note the presence of a memory parameter $M$, which allows sufficiently old measurement/response pairs to be forgotten; this $M$ defines a so-called sliding window mechanism. Such forgetting is critical in a nonstationary bandit problem like that described in Framework 2. Together, $\mathbf{g}_{k}$ and $\mathbf{C}_{k+1}^{-1}$ define the isocontours of a confidence ellipse. The width of the confidence ellipse is parameterized by the upper bound $U_{k}$ on $\|\nabla f(\mathbf{x}_{k})\|$.
With 4, one could take $U_{k}=G$ for all $k$. However, because Algorithm 4 iteratively selects $\mathbf{s}_{k}$ as a maximizer over the confidence ellipse in Algorithm 4, one can imagine that tighter upper bounds $U_{k}$ are generally preferable, so as to encourage better exploitation, that is, better identification of maximizers closer to the mean estimate $\mathbf{g}_{k}$. We will discuss a practical means of dynamically adjusting $U_{k}$ in our numerical experiments. Note, however, that our analysis requires only an upper bound, and so 4 suffices for that purpose.
We set out to prove sublinear regret of Algorithm 4, but we first require two lemmata, the proofs of which are left to the Appendix. The first such lemma bounds the error in the gradient approximation $\mathbf{g}_{k}$ to $\nabla f(\mathbf{x}_{k})$ in arbitrary normalized directions $\mathbf{s}$ in the $k$th iteration of Algorithm 4.
@@
For each $k=1,2,\dots,K$, we have for any $\mathbf{s}\in\mathbb{R}^{d}$ satisfying $\|\mathbf{s}\|=1$,
$\|\mathbf{s}^{\top}(\nabla f(\mathbf{x}_{k})-\mathbf{g}_{k})\|\leq\displaystyle\sqrt{\frac{d(M+1)}{\lambda}}\displaystyle\sum_{i=\ell(k)}^{k-1}\|\nabla f(\mathbf{x}_{i})-\nabla f(\mathbf{x}_{i+1})\|+\sqrt{\lambda}\|\nabla f(\mathbf{x}_{k})\|\|\mathbf{s}\|_{\mathbf{C}_{k}^{-1}},$
where $\ell(k):=\max\{1,k-M-1\}$, and we interpret the summation to be 0 when $k=1$.
The second lemma is a technical result needed to prove Theorem 1; at a high level, Lemma 4 quantifies how the inverse covariance matrix $\mathbf{C}^{-1}$ scales the optimal solution to (5) in each iteration of Algorithm 4. More precisely, Lemma 4 bounds the cumulative sum of these scaled sketches; such results are frequently referred to as elliptical potential lemmata in the linear bandit literature.
@@
Let $M\geq 1$, and denote $\ell(j)=\max\{1,j-M-1\}$. Let $\mathbf{C}_{0}=\lambda\mathbf{I}_{d}$, and let
$\mathbf{C}_{j}=\mathbf{C}_{j-1}+\mathbf{s}_{j}\mathbf{s}_{j}^{\top}-\mathbf{s}_{\ell(j)}\mathbf{s}_{\ell(j)}^{\top}$
for an arbitrary sequence of unit vectors $\{\mathbf{s}_{j}\}_{j=1}^{K}\subset\mathbb{R}^{d}$. Then,
$\sum_{j=1}^{k}\|\mathbf{s}_{j}\|_{\mathbf{C}_{j-1}^{-1}}^{2}\leq\displaystyle\frac{2kd}{M}\log\left(1+\displaystyle\frac{M}{\lambda d}\right).$
We can now prove our main theorem , bounding the dynamic regret of Algorithm 4. To make the statement precise, we define one additional piece of notation, the total variation of the sequence of gradients: that is,
$V_{K}:=\displaystyle\sum_{k=1}^{K}\|\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k})\|.$
@@
Suppose $f$ satisfies 1. Given a sequence of gradients $\{\nabla f(\mathbf{x}_{k})\}$, abbreviate $\mathbf{s}_{k}^{*}:=\nabla f(\mathbf{x}_{k})/\|\nabla f(\mathbf{x}_{k})\|.$ The cumulative dynamic regret of Algorithm 4, with $M\geq 1$, after $K$ iterations is bounded as
$D_{K}\leq\displaystyle\sum_{k=1}^{K}(\mathbf{s}_{k}^{*}-\mathbf{s}_{k})^{\top}\nabla f(\mathbf{x}_{k})\leq 2M\displaystyle\sqrt{\frac{d(M+1)}{\lambda}}V_{K}+\displaystyle\sqrt{\frac{8\lambda d\sum_{k=1}^{K}U_{k}^{2}}{M}\log\left(1+\displaystyle\frac{M}{\lambda d}\right)}\sqrt{K}.$
@@
$\beta_{k}:=\displaystyle\sqrt{\frac{d(M+1)}{\lambda}}\displaystyle\sum_{i=\ell(k)}^{k-1}\|\nabla f(\mathbf{x}_{i+1})-\nabla f(\mathbf{x}_{i})\|.$
@@
$\mathbf{s}_{k}^{*\top}\nabla f(\mathbf{x}_{k})\leq\mathbf{s}_{k}^{*\top}\mathbf{g}_{k}+\beta_{k}+\sqrt{\lambda}\|\nabla f(\mathbf{x}_{k})\|\|\mathbf{s}_{k}^{*}\|_{\mathbf{C}_{k}^{-1}}$
@@
$\mathbf{s}_{k}^{\top}\nabla f(\mathbf{x}_{k})\geq\mathbf{s}_{k}^{\top}\mathbf{g}_{k}-\beta_{k}-\sqrt{\lambda}\|\nabla f(\mathbf{x}_{k})\|\|\mathbf{s}_{k}\|_{\mathbf{C}_{k}^{-1}}.$
@@
$\begin{array}[]{rl}(\mathbf{s}_{k}^{*}-\mathbf{s}_{k})^{\top}\nabla f(\mathbf{x}_{k})&\leq(\mathbf{s}_{k}^{*}-\mathbf{s}_{k})^{\top}\mathbf{g}_{k}+2\beta_{k}+\sqrt{\lambda}\|\nabla f(\mathbf{x}_{k})\|\left[\|\mathbf{s}_{k}^{*}\|_{\mathbf{C}_{k}^{-1}}+\|\mathbf{s}_{k}\|_{\mathbf{C}_{k}^{-1}}\right]\\ &\leq(\mathbf{s}_{k}^{*}-\mathbf{s}_{k})^{\top}\mathbf{g}_{k}+2\beta_{k}+\sqrt{\lambda}U_{k}\left[\|\mathbf{s}_{k}^{*}\|_{\mathbf{C}_{k}^{-1}}+\|\mathbf{s}_{k}\|_{\mathbf{C}_{k}^{-1}}\right]\\ &\leq 2\beta_{k}+2\sqrt{\lambda}U_{k}\|\mathbf{s}_{k}\|_{\mathbf{C}_{k}^{-1}},\end{array}$
where the last inequality is because $\mathbf{s}_{k}$ is a maximizer in eq. 5. Thus,
[…]
where we have used Lemma 4 and the Cauchy–Schwarz inequality to derive the last inequality. ∎
@@ 3.1 Understanding Theorem 1
To better understand Theorem 1, we will focus on Algorithm 2. We first note that, as a very loose bound,
$\|\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k})\|\leq L\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|=L\|\alpha_{k}\mathbf{P}_{k}\nabla f(\mathbf{x}_{k})\|\leq L\|\nabla f(\mathbf{x}_{k})\|$
@@
$V_{K}\leq L\displaystyle\sum_{k=1}^{K}\|\nabla f(\mathbf{x}_{k})\|.$ (8)
We stress that eq. 8 is practically quite loose; for many functions satisfying 1, the bound $\|\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k})\|\leq L\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|$ is not usually tight, given that $L$ is a global Lipschitz constant per 1. Nonetheless, we will use the coarse upper bound in eq. 8 to derive Corollary 1 and Corollary 2. Moreover, we can further upper bound the right-hand side in (8) by using a straightforward result.
@@
Let $K>0$, and let $\delta^{^{\prime}}\in(0,1)$. Let 2 hold, choosing $\delta=\delta^{^{\prime}}/K$, and letting $q\in(0,1)$ be arbitrary. Let 3 hold. Then, Algorithm 2 satisfies, with probability at least $1-\delta^{^{\prime}}$,
$\sum_{k=0}^{K-1}\|\nabla f(\mathbf{x}_{k})\|^{2}\leq\displaystyle\frac{2S_{\max}L(f(\mathbf{x}_{0})-f(\mathbf{x}_{*}))}{q(2\beta-1)}.$ (9)
@@
$\sum_{k=0}^{K-1}\|\nabla f(\mathbf{x}_{k})\|\leq\displaystyle\sqrt{\displaystyle\frac{2S_{\max}L(f(\mathbf{x}_{0})-f(\mathbf{x}_{*}))}{q(2\beta-1)}}\sqrt{K}.$ (10)
@@
By Lemma 1, we have for all $k\leq K-1$
[…]
where 3 was used to derive the last inequality. By 2, with probability at least $1-(\delta^{^{\prime}}/K)$,
$f(\mathbf{x}_{k+1})\leq f(\mathbf{x}_{k})-\frac{q(2\beta-1)}{2S_{\max}L}\|\nabla f(\mathbf{x}_{k})\|^{2}$
@@
[…]
The use of an union bound in the proof of Theorem 2 is admittedly brutish. As a result, for very large horizons $K$, we require the probability $1-(\delta^{^{\prime}}/K)$ in 2 to be so close to 1 that regardless of how $\mathbf{S}_{k}$ is generated, $p_{k}\approx d$. Moreover, Theorem 2 makes sense only when $K$ is finite, which is reasonable for our setting but would be generally nonsensical for performing rate analyses. One route to conducting a more refined analysis is to consider imposing a globalization strategy so that the high-probability convergence rate analysis of [9] could be imported here. Attempting this analysis would complicate our presentation, however, and distract from the purpose of this manuscript. Instead, we interpret Theorem 2 as telling us that, so long as an optimization algorithm used within Framework 2 yields similar first-order optimality guarantees to Algorithm 2, an upper bound like $V_{k}\in\mathcal{O}(\sqrt{K})$ is plausible.
We additionally comment here that the probabilistic nature of Theorem 2 is due entirely to the probabilistic satisfaction of 2. In particular, had we instead made the stronger assumption that (3) is satisfied surely for all $k=1,\dots,K$, then Theorem 2 would also hold surely, since the UCB algorithm behaves deterministically as a function of the input sequence $\{\mathbf{S}_{k}\}$.
We can now state our two corollaries.
@@
Let 4 hold. Let the setting and assumptions of Theorem 2 hold. Suppose Algorithm 2 is employed as the optimization algorithm in Framework 2. Then, with probability at least $1-\delta^{^{\prime}}$,
$D_{K}\leq\left(2M\displaystyle\sqrt{\frac{2d(M+1)S_{\max}L(f(\mathbf{x}_{0})-f(\mathbf{x}_{*}))}{q(2\beta-1)\lambda}}+\displaystyle\sqrt{\frac{8\lambda dKG^{2}}{M}\log\left(1+\displaystyle\frac{M}{\lambda d}\right)}\right)\sqrt{K},$
and so if we choose $M=\lceil K^{\frac{1}{4}}\rceil$, where $\lceil\cdot\rceil$ denotes the ceiling function, then
@@
Additionally, we can prove a result for the idealized case where $U_{k}=\|\nabla f(\mathbf{x}_{k})\|$ (that is, the upper bound used in eq. 5 is maximally tight) on every iteration.